OpenStudy (anonymous):

PLEASE HELP f(x)=log(x+5)-3 a) domain of f is______ b) graph c) range of f is___ vertical asymptote is___ or No asymptote d) Find f^-1, the inverse of f e) domain of f^-1 range of f^-1 f) Graph f^-1

OpenStudy (anonymous):

To start are there any you think you know how to do already?

OpenStudy (anonymous):

I know that when I find the domain and range of f(x) it would be the opposite for f^-1(x) I know the graph already but I just dont know how to find the domain and range of vertical asymptote

OpenStudy (anonymous):

Okay, for logarithms I believe that the range will be (-INFINITY, INFINITY) every time. The domain can be found by knowing where a logarithmic function DNE. This would be at log(0) so to find domain, just set the inside equal to 0 and solve for x. Your domain would then be (x, INFINITY). Then you can almost guess the vertical asymptote from there because since it will head off to negative infinity, but never reach x, you know that that must be the vertical asymptote.

OpenStudy (anonymous):

so the range is (-infinity, infinity) the vertical asymptote is (-2), how would you do the domain? still confused

OpenStudy (anonymous):

Careful. Asymptote wouldn't be (-2) because the 3 is outside of the log function. So all that is doing is shifting the graph down on the y-axis. As for domain, log(x) exists for ALL x > 0. So we have log(x+5) and we know that (x+5) > 0. Solve for x and we find that x > -5, therefore the domain is (-5, INFINITY).

OpenStudy (anonymous):

so the vertical asymptote would be -5?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

how would you get the inverse?

OpenStudy (anonymous):

of the equation?

OpenStudy (anonymous):

Okay, so here we need to use what we know about inverses.\[f(f^{-1}(x)) = x\]For us, \[f(x) = log(x+5)-3\]so then, combine those two equations and we find that \[f(f^{-1}(x))=log(f^{-1}(x)+5)-3=x\]From here we would use properties of logs to find \[f^{-1}(x)\]So remember that \[\log_{b}(a)=x \] tells us that \[b^{x} = a\]

OpenStudy (anonymous):

okay then how would it look like?

OpenStudy (anonymous):

Okay, so we plugged in \[f^{-1}(x)\] for x in f(x). Whenever you're given just log without a base it is really \[log_{10}\]. So we know the equation is equal to x. We want the log by itself on one side so we can apply the basic property of logarithms so\[log_{10}(f^{-1}(x) +5) -3 =x\] becomes\[log_{10}(f^{-1}(x) +5) = x+3\]Apply the property of logs we know and we find that \[10^{x+3} = f^{-1}(x) +5\] subtract 5 from each side, and we see that \[f^{-1}(x) = 10^{x+3} - 5\]

OpenStudy (anonymous):

okay i can see how you got the answer now! thank you super much!!!

OpenStudy (anonymous):

Your welcome! :)

OpenStudy (anonymous):

can you help me find the inverse of (5/(x-3))

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Okay, the key to finding pretty much any inverse function is knowing that \[f(f^{-1}(x))=x\]Suppose that you have the function \[f(x) = x+2\]and you want to find \[f(5)\]How would you do that?

OpenStudy (anonymous):

plug in 5 for x

OpenStudy (anonymous):

Exactly! And it is precisely the same approach that we will use to find the inverse. If \[f(f^{-1}(x)) =x\] and \[f(x)=\frac{5}{x-3}\]then how could we express \[f(f^{-1}(x))?\]

OpenStudy (anonymous):

so it will look like\[\frac{ 5 }{ f ^{-1} -3}\]

OpenStudy (anonymous):

to start out with first

OpenStudy (anonymous):

Yes. And then that would be equal to what?

OpenStudy (anonymous):

\[f ^{-1}=5+3 ???\]

OpenStudy (anonymous):

not quite. Remember that \[f(f^{1}(x)) = x\] so when we plugged in \[f^{-1} (x)\] to the function, we know that it will be equal to x. So then we have \[\frac{5}{f^{-1}(x)-3}=x\]Leaving us with just needing to solve for \[f^{-1}(x)\]

OpenStudy (anonymous):

so it would be \[f ^{-1}=5x+3\]

OpenStudy (anonymous):

Close. Remember that when you have a plus or minus sign with the x you need to be careful how you reorder terms. We would want to treat the bottom as though it were all in parenthesis like so\[\frac{5}{(f^{-1}(x)-3)}=x\]This makes the algebra a bit messy. The way I find best to keep the terms straight is to multiply the denominator by both sides so that you get\[5=x(f^{-1}(x)-3)\]then divide by x giving you \[\frac{5}{x}=f^{-1}(x)-3\]and finally add 3 to both sides to get \[\frac{5}{x}-3=f^{-1}(x)\]

OpenStudy (anonymous):

The other way you could solve it is by dividing both sides by 5 first so that you have \[\frac{1}{f^{-1}(x) - 3}=\frac{x}{5}\]And then you invert both fractions so that you have \[f^{-1}(x)-3=\frac{5}{x}\]And finally subtract 3\[f^{-1}(x)=\frac{5}{x}-3\]The reason why I prefer to do it the other way is I feel that when you do it this way you become prone to forgetting to invert both sides so you'll only flip the side you are trying to solve for thereby changing the actual equation.

OpenStudy (anonymous):

okay that makes way more sense when i plug in the calculator

OpenStudy (anonymous):

okay thank you sooooooooo MUCH!!!!!!!!!!!!!!!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

but wouldn't it be +3

OpenStudy (anonymous):

oops. yes haha. good catch.

OpenStudy (anonymous):

thank you again!